OpenStudy (idani):

Factor completely x 3 +2x 2 ­-15x A. x(x-3)(x+5) B. -3x(x+5)(x+1) C. 5x(x+1)(x-3) D. x(x+5)(x+3)

jhonyy9 (jhonyy9):

x^3 +2x^2 -15x = x(x^2 +2x -15) = ? do you can continue it ?

jhonyy9 (jhonyy9):

for this inside parentheses you need getting two integers with sum = 2 and product equal 15 hope this will help you

jhonyy9 (jhonyy9):

oh sorry product equal minus 15

jhonyy9 (jhonyy9):

x^2 +2x -15 = x*x +5x -3x +5*(-3) = x(x+5) -3(x+5) = (x-3)(x+5) ok. ?

jhonyy9 (jhonyy9):

sorry in the front pf parentheses i missed the x

jhonyy9 (jhonyy9):

so will be x(x-3)(x+5)

jhonyy9 (jhonyy9):

so choice A. is right sure

jhonyy9 (jhonyy9):

do you understand it now sure ?

OpenStudy (idani):

Yes thank you so much @jhonyy9

jhonyy9 (jhonyy9):

was my pleasure and anytime good luck bye bye